How long will it take a drop of dextrose from a bottle 2.30 m high hit the floor?


1
Expert's answer
2021-11-19T10:50:37-0500

We can find the time that the drop of dextrose takes to fall from the height of 2.30 meters as follows:


"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\times2.30\ m}{9.8\ \dfrac{m}{s^2}}}=0.68\ s."